Matthews & Bros. is a local landscape construction company. In analyzing financial performance, the cost accountant compares actual results with a flexible budget. The standard direct labor rates used in the flexible budget are established each year at the time the annual plan is formulated and held constant for the entire year.

The standard direct labor rates in effect for the current fiscal year and the standard hours allowed for the actual output of work for July are shown in the following schedule.

Worker Classification Standard Direct Labor Rate per Hour Standard Direct Labor-Hours Allowed for Output
I $ 40 1,500
II 35 2,000
III 25 2,500

The wage rates for each labor class increased under the terms of a new contract. The standard wage rates were not revised to reflect the new contract.

The actual direct labor-hours worked and the actual direct labor rates per hour experienced for the month of July were as follows.

Worker Classification Actual Direct Labor Rate per Hour Actual Direct Labor-Hours
I $ 44 1,660
II 38 2,600
III 26 1,800

Required:

Calculate the dollar amount of the total direct labor variance for July for Matthews & Bros. and break down the total variance into the following components:

a. Direct labor price and efficiency variances.

b. Direct labor mix and yield variances.
Answer #1
Direct Labor Price Variance
Formula : Actual Hours Worked * ( Standard Price per hour - actual price per hour)
Worker Classification Actual Hours Worked Standard Price Actual Price Variance
a b c d e=b(c-d)
I 1660 40 44 6640 Unfavorable
II 2600 35 38 7800 Unfavorable
III 1800 25 26 1800 Unfavorable
Net 16240 Unfavorable
Direct Labor Efficiency Variance
Formula :Standard Rate Per Hour * (Standard Hours - Actual Hours)
Worker Classification Standard Price Standard Hours Actual Hours Worked Variance
a b c d e=b*(c-d)
I 40 1500 1660 6400 Unfavorable
II 35 2000 2600 21000 Unfavorable
III 25 2500 1800 17500 Favorable
Net 6000 6060 9900 Unfavorable
Direct Labour Mix Variance
= Standard Rate * (Standard Hours - Actual Hours In Standard Mix)
Worker Classification Standard Price Standard Hours Actual Hours In Std mix Variance
a b c d e=b*(c-d)
I 40 1500 1515 600 Unfavorable
II 35 2000 2020 700 Unfavorable
III 25 2500 2525 625 Unfavorable
Net 6000 6060 1925 Unfavorable
Direct Labor Yeild Variance
= Standard Rate * (Actual Hours In Standard Mix - Actual Hours)
Worker Classification Standard Price Actual Hours In Std mix Actual Hours Worked Variance
a b c d e=b*(c-d)
I 40 1515 1660 5800 Unfavorable
II 35 2020 2600 20300 Unfavorable
III 25 2525 1800 18125 Favorable
Net 6060 6060 7975 Unfavorable